Consider the following four--‐step assembly operation with quality problems. -The first resource has a processing time of 5 minutes per unit and one employee doing the operation ‐ The second resource has a processing time of 4 minutes per unit. It also has one employee doing the operation. However, this is a very delicate task and 0.85 of all products have to be scrapped after this step. ‐ Two workers are staffed for the third resource. No quality problems occur at this resource and the processing time is 22minutes per unit. ‐ At the fourth and final resource, one operator handles the product. No quality problems exist at this step and the processing time is 12 minutes per unit. What is the process capacity per hour? (keep two decimal)
Answer #1

For process capability, we can ignore the number of resources and consider the workstations here.

Step 1: Processing time 5 minute per unit. The capacity = 60/5 = 12 units per hour.

Step 2: Processing time 4 minutes per unit. The capacity = 60/4 = 15 units. Out of this 0.85 of them are scrapped. Thus net capacity is 15*(1-0.85) = 2.25 units per hour.

Step 3: Processing time 22 minutes per unit. The capacity = 60/22 = 2.72 units per hour.

Step 4: Processing time 12 minutes per unit. The capacity = 60/12 = 5 units per hour.

Out of these processes. The bottleneck is created at step 2 and that becomes the overall process capacity per hour.

The answer: 2.25 units per hour.